OpenStudy (anonymous):

SolutionS of x^2+12x+10=0

OpenStudy (anonymous):

-1 and -11

OpenStudy (accessdenied):

How do you get -1 and -11?

OpenStudy (anonymous):

Is it right?

OpenStudy (accessdenied):

No, since x=-1 and x=-11 being solutions =>x+1=0 and x+11=0 (x+1)(x+11) = x^2 + 11x + x + 11 = x^2 + 12x + 11, not x^2 + 12x + 10

OpenStudy (anonymous):

I don't know then

OpenStudy (accessdenied):

The discriminant is 104, which is not a perfect square and thus indicates that there are two solutions that are irrational. I'll use quadratic formula since I already have the discriminant: \[ x = \frac{-b \pm \sqrt{D}}{2a} \] \[ x=\frac{-12 \pm \sqrt{104}}{2(1)}\\ = \frac{-12 \pm \sqrt{4\times26}}{2}\\ = \frac{-\cancel{12}6 \pm \cancel2\sqrt{26}}{\cancel2}\\ x= -6 \pm \sqrt{26} \]

OpenStudy (anonymous):

That's what I got but I just take the real number

OpenStudy (accessdenied):

I see now, though it would be more accurate to use \(x=-6 \pm \sqrt{26}\) than approximating to nearest integer :)

OpenStudy (anonymous):

yea ok

OpenStudy (anonymous):

my bad